OpenStudy (anonymous):

Find the median for the following set of data. 8, 3, 7, 4, 7, 4, 2, 5 4 5 4.5 5.5

OpenStudy (katrinakaif):

5.5 Add 7 + 4 = 11 ad divide by 2

OpenStudy (anonymous):

2,3,4,4,5,7,7,8 9/2=4.5

OpenStudy (anonymous):

4.5

OpenStudy (anonymous):

you must put those numbers in order

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

(4+5)/2=9/2=4.5

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw :)